The present value of a series of payments of $500 at the beginning of every month years, forever, is equal to $10000. Calculate the nominal annual discount rate compounded monthly.
1
Expert's answer
2020-09-28T20:11:37-0400

solution


"payment, \ p = 500" is paid "monthly\ in \ perpetuity" and "present\ value ,\ pv =10000"


These represents a series of payments beginning at time 0, i.e annuity in advance


"present\ value =p*\frac{1+i}{i}"


Where "i" is the monthly compounding rate


"10000 =500*\frac{1+i}{i}""20=\frac{1+i}{i}""i=\frac{1}{19}=5.2632\%"

Answer

The nominal annual discount rate:


"(5.2632\% )*12=63.1579\%"